# Log-Scale Quantization in Distributed First-Order Methods: Gradient-based Learning from Distributed Data

### v2

@@ Abstract
Decentralized strategies are of interest for learning from large-scale data over networks. This paper studies learning over a network of geographically distributed nodes/agents subject to quantization. Each node possesses a private local cost function, collectively contributing to a global cost function, which the proposed methodology aims to minimize. In contrast to many existing literature, the information exchange among nodes is quantized. We adopt a first-order computationally-efficient distributed optimization algorithm (with no extra inner consensus loop) that leverages node-level gradient correction based on local data and network-level gradient aggregation only over nearby nodes. This method only requires balanced networks with no need for stochastic weight design. It can handle log-scale quantized data exchange over possibly time-varying and switching network setups. We analyze convergence over both structured networks (for example, training over data-centers) and ad-hoc multi-agent networks (for example, training over dynamic robotic networks). Through analysis and experimental validation, we show that (i) structured networks generally result in a smaller optimality gap, and (ii) logarithmic quantization leads to smaller optimality gap compared to uniform quantization.
@@
Motivated by recent developments in cloud-computing and parallel processing along with low-cost CPUs and communication networks, this paper considers distributed/decentralized algorithms for machine learning and optimization. This work finds a wide range of practical applications in decentralized data mining, where the data set is distributed among a network of computing nodes. In the real-world scenarios, practical model nonlinearities such as data quantization must be addressed for information-exchange among the computing nodes. The primary emphasis of this work is to address log-scale quantization and compare it versus uniform quantization. Further, the impact of the structure of the data mining network is studied to reduce the optimality gap and improve the convergence rate. As contemporary distributed and networked data mining systems demand for more accurate algorithms with fast rate of convergence for real-time applications, our research highlights the intricate use of logarithmic quantization over structured networks. In particular, our results can be extended to other machine learning algorithms for more accurate and faster data mining solutions.
@@ I Introduction
Learning from geographically distributed data has emerged as a critical challenge in various domains such as sensor networks, edge computing, and decentralized systems [1, 2, 3]. The scenario involves individual nodes possessing private local cost functions, collectively contributing to a global cost that needs to be minimized [4, 5]. This paper focuses on addressing this challenge through the lens of gradient-based optimization over networks, targeting scenarios where the cost functions are locally known at distributed nodes communicating over quantized channels. Data quantization, the process of representing data with a reduced number of bits, is often employed to reduce the communication load in such decentralized setups and is proven as an effective way to reduce communication costs, for example, in machine learning applications [6, 7]. One common type of quantization is logarithmic or log-scale that efficiently assigns more bits to represent smaller values and fewer bits to larger values [8, 9]. This is beneficial when dealing with gradients or weights that often have a wide dynamic range. Smaller values, which might be more critical for convergence, are represented with higher precision. This may increase the complexity of the distributed learning algorithm as we move from linear dynamics to non-linear dynamics due to quantization. Uniform quantization, on the other hand, simplifies the representation of values, which can lead to a faster communication process. However, the lack of precision, especially for small values, might affect performance in scenarios where accurate representation of gradients or weights is crucial, and lead to residual error [10, 11]. In this paper, we study both uniform and log-scale quantization for distributed optimization.
@@
The existing literature on distributed optimization is mostly linear, where the focus is mainly on reducing the optimality gap and improving the convergence rate. The primary distributed optimization methods are GP [12] and DGD algorithm [13], which are further extended to DSGD [14] and SGP [15, 16] by adopting stochastic gradient methods. Algorithm ADDOPT [17] and its stochastic version S-ADDOPT [18] further accelerate the convergence rate using gradient tracking. Similarly, momentum-based approaches improve the convergence rate using heavy-ball method [19], Nesterov gradient method [20, 21], or a combination of both [22]. On the other hand, PushSVRG [23] and PushSAGA [24, 25] algorithms use variance reduction to eliminate the uncertainty caused by the stochastic gradients. Few other distributed optimization works use sign-based nonlinearity in the node dynamics to address finite-time and fixed-time convergence [26, 27, 28]. All these methods assume unlimited bandwidth to communicate arbitrary real numbers which in practice is not possible as all realistic communication protocols rely on finite precision. The effect of uniform quantization over channels with limited data-rate and the associated optimality gap (or residual) is discussed in [29, 30, 31]. What is missing from the existing literature is the consideration of log-scale quantization for distributed optimization. This is primarily addressed for distributed resource allocation and constraint-coupled optimization [32], but not for distributed learning over networks.
@@ II Assumptions, Terminology, and Problem Statement
where $\rho$ is the uniform quantization level. The two quantization functions are represented in Fig. 1 for comparison. It can be observed that, in contrast to the log-scale quantization, the upper/lower bounds of uniform quantization, defined as $z\pm\dfrac{\rho}{2}$, are biased from the origin. In this work, we study the convergence under log-scale quantization as compared to uniform quantization. These two are the main types of the nonlinearity governing the data-transmission setup.
@@
Each local cost function $f_{i}$ has $L$-smooth and is not necessarily convex, while the global cost $F$ is strictly convex.
These standard assumptions imply that the global cost function is differentiable and its second derivative is bounded above by $L$ and below by $\mu$. Further, the global optimizer $\mathbf{x}^{*}$ of the cost $F$ exists and is also unique. The followig subsection further motivates the study in this paper.
@@ II-A Effect of Network Structure and Quantization
This paper investigates the effect of network structure by numerical experimentation. In distributed optimization and learning over networks, the choice of network topology plays a crucial role in determining the convergence rate and optimality gap of the learning algorithms [36, 37]. Different network topologies can have varying effects on the efficiency and speed of convergence. A general experimental comparison between structured networks (such as exponential graphs [^2] In exponential graphs, as the number of nodes increases exponentially, the degree at each node increases linearly.) and ad-hoc networking setups (such as geometric graphs [^3] In geometric graphs, two nodes are connected if they are in physical proximity.) is considered in this paper. Examples of structured networks include networked data-centers or hierarchical social networks, and examples of ad-hoc random networks include collaborative robotic networks or wireless sensor networks. Structured networks (such as exponential graphs), characterized by well-defined and organized connections, can often lead to faster convergence rates. In exponential graphs, where nodes have a hierarchical structure, information can propagate more efficiently through the network. Furthermore, structured networks may exhibit lower optimality gaps due to better coordination among the nodes. The organized nature of the network facilitates synchronized updates, potentially leading to quicker convergence to a near-optimal solution. On the other hand, ad-hoc multi-agent networks (such as random geometric graphs), which might lack a clear structure, can have a more complex convergence behaviour. The convergence rate may be slower with a higher optimality gap compared to structured networks because information propagation might not follow an organized pattern. Nodes may require more iterations to align their models and achieve a consensus on the optimal solution.
@@
Another concern is to consider switching network topologies addressing time-varying setups, for example, in mobile sensor networks and collaborative swarm robotics. The change in the network topology might be due to packet drops or link failures. Recall that most existing linear algorithms [12, 13, 14, 15, 16, 17, 18, 19, 21, 22, 23, 24] necessitate stochastic weight design which is prone to change in the network topology. This is because any change in the network structure violates the stochastic condition and requires the redesign of stochastic weights, for example by applying the algorithms in [38, 39, 40]. In contrast, this work only requires WB design of matrices $A_{\gamma}$ and $B_{\gamma}$ that allows handling possible changes in the network topology. This is better illustrated in Fig. 2.
@@
We address more findings in terms of data quantization over channels. First, it should be noted that logarithmic quantization, as compared to uniform quantization, is more efficient in scenarios where the optimization problem involves a wide range of gradient magnitudes. This follows the sector-bound nature of the log-scale quantization, as illustrated in Fig. 1. Smaller values, which are key for convergence, can be represented more accurately with logarithmic quantization. This may lead to a smaller optimality gap and more efficient convergence. On the other hand, the lack of precision in uniform quantization, especially for small values, might slow down convergence and increase the optimality gap in the range of values where accurate representation of gradients or weights is crucial. In terms of network structure in the presence of quantization, although exponential graphs are structured and can facilitate efficient quantized communication, their optimality gap might be considerably affected by quantization inaccuracy. In other words, the efficacy of the network structure might be dominated by the effect of quantization. These comments are investigated by the numerical experiments in Section IV.
@@ III-A Algorithm
This section presents our main algorithm for learning over networks. The methodology is based on two WB matrices $A_{\gamma}=[a^{\gamma}_{ij}]\in\mathbb{R}^{n\times n}$ and $B_{\gamma}=[b^{\gamma}_{ij}]\in\mathbb{R}^{n\times n}$ that capture the weights for data-sharing over the network, one for consensus on the states and the other for gradient-tracking dynamics. The index $\gamma:k\mapsto\Gamma$ denotes the associated network topology that may change over time, whereas $\Gamma$ denotes the set of all possible topologies of the data-sharing network. The graph is all-time connected, i.e., all the netwotk topologies in $\Gamma$ are connected. The two matrices $A_{\gamma}$ and $B_{\gamma}$ can be designed independently, matrix $A_{\gamma}$ to characterize the weights on the local state value $\mathbf{x}_{i}$ and matrix $B_{\gamma}$ to characterize the weights on the auxiliary variable $\mathbf{y}_{i}$ for gradient tracking. These weight matrices are WB over $\mathcal{G}$ as discussed before, where the entries satisfy the following:
@@
In contrast to most existing methods with weight-stochastic matrices (introduced in Section I), the proposed dynamics work under WB design and under switching network topologies. These make it easier to handle link-failure as discussed in Fig. 2. Moreover, the link (or channel) nonlinearity $h_{l}(\cdot)$ is not considered in the existing literature, which allows to address log-scale data quantization by setting $h_{l}(z)=q_{l}(z)$ given in Eq. (4). The discrete-time version of our methodology is summarized in Algorithm 1.
@@ III-B Proof of Convergence
This section proves the convergence of the distributed dynamics (19)-(20) using eigen-spectrum perturbation-based analysis. We also state some relevant lemmas. The proofs in this section hold for any sign-preserving, odd, sector-bound nonlinear mapping $h_{l}(\cdot)$, wich also includes log-scale quantization $h_{l}(z)=q_{l}(z)$ defined by (4). In the rest of the paper, for notation simplicity, we drop the dependence $(t,\gamma)$ unless where it is needed.
@@
[43] Given the balanced digraph $\mathcal{G}$/ its Laplacian matrix $\overline{A}$ (or $\overline{B}$) has all its eigenvalue in the left-half-plane except one isolated zero eigenvalue with associated non-negative left eigenvector $\mathbf{u}_{1}^{\top}$ satisfying $\sum_{i=1}^{n}u_{1,i}>0$.
@@
where $M(\alpha)=M_{0}+\alpha M_{1}$ with
@@
For the rest of the proof, we study the eigenspectrum perturbation of $M^{0}$ and then extend the results to $M^{0}_{q}$ using Eq. (44). From Lemma 2, both Laplacian matrices  $\overline{A}$ and $\overline{B}$ have one isolated zero eigenvalue and the rest in the left-half-plane. Therefore, $m$ sets of eigenvalues of $M_{0}$, associated with dimensions $j=\{1,\ldots,m\}$ are in the form,
@@
Next, we consider $\alpha M_{1}$ as a perturbation to matrix $M_{0}$. This is done using Lemma 1. In particular, we check how the perturbation term $\alpha M_{1}$ affects the two zero eigenvalues $\lambda_{1,j}$ and $\lambda_{2,j}$ of $M_{0}$. Denote the perturbed zero eigenvalues by $\lambda_{1,j}(\alpha)$ and $\lambda_{2,j}(\alpha)$ associated to $M$. The right and left unit eigenvectors of $\lambda_{1,j}$$\lambda_{2,j}$ follow from Lemma 2 and [43] as [^5] The normalizing factors of the unit vectors might be ignored as in the followings we only care about the sign of the terms in $U^{\top}M_{1}V$, not the exact values. ,
@@
where $u_{2,i}$ is the $i$th element of $\mathbf{u}_{2}$. Then, using Lemma 1, the perturbations  $\partial_{\alpha}d\lambda_{1,j}|_{\alpha=0}$ and $\partial_{\alpha}d\lambda_{2,j}|_{\alpha=0}$ are defined based on the eigenvalues of the matrix in (54). Since this matrix is triangular and from Eq. (55), $\partial_{\alpha}\lambda_{1,j}|_{\alpha=0}=0$ and $\partial_{\alpha}\lambda_{2,j}|_{\alpha=0}<0$. Therefore, from the perturbation analysis, one can conclude that $\alpha M_{1}$ pushes $m$ zero eigenvalues $\lambda_{2,j}(\alpha)$ of $M$ toward the left-half-plane and the other $m$ zero eigenvalues $\lambda_{1,j}(\alpha)$ remain at zero. This implies that sufficiently small $\alpha$ gives the eigen-spectrum of $M$ as
@@
First, we determine the admissible bound on $\alpha$ for proof of convergence. Recall from [45, Appendix] that one can relate the spectrum of $M(\alpha)$ in (31) to $\alpha$. For ease of notation, from this point onward we derive the bound for $m=1$ (but it holds for any $m>1$). Performing row/column permutations in [45, Eq. (18)] $\sigma(M)$ can be determined from the following,
@@
Similar to the proof of Theorem 1, for stability we need to find the admissible range of $\alpha$ values for which the eigenvalues $\lambda$ remain in the left-half-plane, except one zero eigenvalue. The analysis here is based on the fact that the eigenvalues are continuous functions of the matrix elements [46]. It is clear that $\alpha=0$ satisfies Eq. (57) and gives $\mbox{det}((\overline{W}-\lambda I_{n})(\overline{A}-\lambda I_{n}))=0$. This leads to the eigen-spectrum following as $\sigma(M)=\sigma(\overline{A})\cup\sigma(\overline{W})$ with two zero eigenvalues. We need to find the other root $\overline{\alpha}>0$, which gives the admissible range as $0<\alpha<\overline{\alpha}$ for the stability of $M(\alpha)$. This follows the continuity of $\sigma(M)$ as a function of $\alpha$ [46]. With some abuse of notation, for any $\operatorname{Re}\{\lambda\}<0$ one can reformulate (57) as
@@
This implies that $\dot{\mathcal{V}}={\delta}^{\top}\dot{\delta}=\delta^{\top}M_{q}{\delta}$ and from [43, Sections VIII-IX], we have

### v3

@@ Abstract
Decentralized strategies are of interest for learning from large-scale data over networks. This paper studies learning over a network of geographically distributed nodes/agents subject to quantization. Each node possesses a private local cost function, collectively contributing to a global cost function, which the considered methodology aims to minimize. In contrast to many existing papers, the information exchange among nodes is log-quantized to address limited network-bandwidth in practical situations. We consider a first-order computationally efficient distributed optimization algorithm (with no extra inner consensus loop) that leverages node-level gradient correction based on local data and network-level gradient aggregation only over nearby nodes. This method only requires balanced networks with no need for stochastic weight design. It can handle log-scale quantized data exchange over possibly time-varying and switching network setups. We study convergence over both structured networks (for example, training over data-centers) and ad-hoc multi-agent networks (for example, training over dynamic robotic networks). Through experimental validation, we show that (i) structured networks generally result in a smaller optimality gap, and (ii) log-scale quantization leads to a smaller optimality gap compared to uniform quantization.
@@
Motivated by recent developments in cloud computing, parallel processing, and the availability of low-cost CPUs and communication networks, this paper considers distributed and decentralized algorithms for machine learning and optimization. These algorithms are particularly relevant for decentralized data mining, where data sets are distributed across a network of computing nodes. A practical example of this is the classification of images over a networked data centre. In real-world scenarios, practical model nonlinearities such as data quantization must be addressed for information exchange among the computing nodes. This work emphasizes the importance of handling log-scale quantization and compares its performance over uniform quantization. By exploring these quantization methods, we aim to determine which is more accurate in terms of optimality gap and learning residual. Moreover, we study the impact of the structure of the information-sharing network on reducing the optimality gap and improving the convergence rate of distributed algorithms. As contemporary distributed and networked data mining systems demand highly accurate algorithms with fast convergence for real-time applications, our research emphasizes the benefit of structured networks under logarithmic quantization information-exchange. Our findings can be extended to different machine learning algorithms, offering pathways to more accurate and faster data mining solutions.
@@ I Introduction
Learning from geographically distributed data has emerged as a critical challenge in various domains such as sensor networks, edge computing, and decentralized systems [1, 2, 3]. The scenario involves individual nodes possessing private local cost functions, collectively contributing to a global cost that needs to be minimized [4, 5, 6, 7]. This paper focuses on addressing this challenge through gradient-based optimization over networks, targeting scenarios where the cost functions are locally known at distributed nodes communicating over quantized channels. Data quantization, the process of representing data with a reduced number of bits, is often employed to reduce the communication load in such decentralized setups and is proven as an effective way to reduce communication costs, for example, in machine learning applications [8, 9]. One common quantisation type is logarithmic or log-scale that efficiently assigns more bits to represent smaller values and fewer bits to larger values [10, 11]. This is beneficial when dealing with gradients or weights that often have a wide dynamic range. Smaller values, which might be more critical for convergence, are represented with higher precision. This may increase the complexity of the distributed learning algorithm as we move from linear dynamics to non-linear dynamics due to quantization. Uniform quantization, on the other hand, simplifies the representation of values, which can lead to a faster communication process. However, the lack of precision, especially for small values, might affect performance in scenarios where accurate representation of gradients or weights is crucial, and lead to residual error [12, 13]. Therefore, studying and comparing uniform and log-scale quantization for distributed optimization is of interest for distributed data mining applications.
@@
The existing literature on distributed optimization mostly consider linear function of the state variables, where the focus is mainly on reducing the optimality gap and improving the convergence rate. The primary distributed optimization methods are GP [14] and DGD algorithm [15], which are further extended to DSGD [16] and SGP [17, 18] by adopting stochastic gradient methods. Algorithm ADDOPT [19] and its stochastic version S-ADDOPT [20] further accelerate the convergence rate using gradient tracking. Similarly, momentum-based approaches improve the convergence rate using heavy-ball method [21], Nesterov gradient method [22, 23], or a combination of both [24]. On the other hand, PushSVRG [25] and PushSAGA [26, 27] algorithms use variance reduction to eliminate the uncertainty caused by the stochastic gradients. Few other distributed optimization works use sign-based nonlinearity in the node dynamics to address finite-time and fixed-time convergence [28, 29, 30]. All these methods assume unlimited bandwidth to communicate arbitrary real numbers which in practice is not possible as all realistic communication protocols rely on finite precision. The effect of uniform quantization over channels with limited data-rate and the associated optimality gap (or residual) is discussed in [31, 32, 33]. The work [34] proposes a novel quantized distributed gradient tracking algorithm to minimize the finite sum of local costs over digraphs with linear convergence rate. The paper [35] combines the gradient tracking Push-Pull method with communication compression (quantized links). Similarly, communication compression technique compatible with a general class of operators unifying both unbiased and biased compressors is discussed in [36]. Distributed quantized gradient-tracking under a dynamic encoding and decoding scheme is considered in [37]. What is missing from the existing literature is the consideration of log-scale quantization for distributed optimization to address resource-constrained communication networks with limited-bandwidth. This is primarily addressed for distributed resource allocation and constraint-coupled optimization [38], but not for distributed learning over networks.
@@ II Assumptions, Terminology, and Problem Statement
where $\rho$ is the uniform quantization level. The two quantization functions are represented in Fig. 1 for comparison. It can be observed that, in contrast to the log-scale quantization, the upper/lower bounds of uniform quantization, defined as $z\pm\dfrac{\rho}{2}$, are biased from the origin. In this work, we study the convergence under log-scale quantization as compared to uniform quantization. These two are the main types of the nonlinearities governing the data-transmission setup; for example see the following references on uniform quantization [34, 35, 36, 37] and on logarithmic quantization [42, 38]. These nonlinearities are employed to reduce the communication bandwidth and network traffic in resource-constrained networking setups.
@@
The WB condition implies that the sum of the weights on incoming links and outgoing links are equal at all nodes. This holds for every time-instant and for all the configurations of the network topology. To give a motivation behind this assumption, note that the WB condition is more relaxed than weight-stochasticity assumption in related works which require the link weights to sum to one for all nodes.
@@
Each local cost function $f_{i}$ has $L$-smooth and is not necessarily convex, while the global cost $F$ is strongly convex.
The above assumption is motivated by the discussion in [43] and is used in many references therein saying that it is common to use strong convexity of $F$ as many practical optimization objectives tend to be strongly convex near the minimum. These standard assumptions imply that the global cost function is differentiable and its second derivative is bounded above by some constant $L$ and below by strong-convexity constant $\mu$ where $0<\mu<L$. Further, the global optimizer $\mathbf{x}^{*}$ of the cost $F$ exists and is also unique. The following subsection further motivates the study in this paper.
@@ II-A Effect of Network Structure and Quantization
This paper investigates the effect of network structure by numerical experimentation. In distributed optimization and learning over networks, the choice of network topology plays a crucial role in determining the convergence rate and optimality gap of the learning algorithms [44, 45]. Different network topologies can have varying effects on the efficiency and speed of convergence [46, 47, 48]. A general experimental comparison between structured networks (such as exponential graphs [^2] In exponential graphs, as the number of nodes increases exponentially, the degree at each node increases linearly.) and ad-hoc networking setups (such as geometric graphs [^3] In geometric graphs, two nodes are connected if they are in physical proximity.) is considered in this paper. Examples of structured networks include networked data centres or hierarchical social networks, and examples of ad-hoc random networks include collaborative robotic networks or wireless sensor networks. Structured networks (such as exponential graphs), characterized by well-defined and organized connections, can often lead to faster convergence rates [45]. In exponential graphs, where nodes have a hierarchical structure, information can propagate more efficiently through the network [49]. Furthermore, structured networks may exhibit lower optimality gaps due to better coordination among the nodes [44]. The organized nature of the network facilitates synchronized updates, potentially leading to quicker convergence to a near-optimal solution. On the other hand, ad-hoc multi-agent networks (such as random geometric graphs), which might lack a clear structure, can have a more complex convergence behaviour. The convergence rate may be slower with a higher optimality gap compared to structured networks because information propagation might not follow an organized pattern. Nodes may require more iterations to align their models and achieve a consensus on the optimal solution [50].
@@
Another concern is to consider switching network topologies addressing time-varying setups, for example, in mobile sensor networks and collaborative swarm robotics. The change in the network topology might be due to packet drops or link failures. Recall that most existing linear algorithms [14, 15, 16, 17, 18, 19, 20, 21, 23, 24, 25, 26] necessitate stochastic weight design which is prone to change in the network topology. This is because any change in the network structure violates the stochastic condition and requires the redesign of stochastic weights, for example by applying the algorithms in [51, 52, 53]. In contrast, this work only requires WB design of matrices $A_{\gamma}$ and $B_{\gamma}$ that allows handling possible changes in the network topology. This is better illustrated in Fig. 2. To preserve the WB condition in the presence of packet drops, we assume that if a packet is dropped on $i\rightarrow j$ link then $j\rightarrow i$ link is automatically deleted or, in other words, the agent $i$ does not apply the packet received from $j$ (for example, by setting the associated weight equal to zero), and the entire link is removed. This assumption is justified in existing literature [52, 54] saying that assuming common knowledge, both agents $i,j$ are aware of the delivery or loss of the packets over the mutual link.
@@
We address more findings in terms of data quantization over channels. First, it should be noted that logarithmic quantization, as compared to uniform quantization, is more efficient in scenarios where the optimization problem involves a wide range of gradient magnitudes. This follows the sector-bound nature of the log-scale quantization, as illustrated in Fig. 1. Smaller values, which are key for convergence, can be represented more accurately (with more bits) under logarithmic quantization. This may lead to a smaller optimality gap and more efficient convergence, for example in optimal resource allocation [55]. The lack of precision in uniform quantization, especially for small values, might slow down convergence and increase the optimality gap in the range of values where accurate representation of gradients or weights is crucial [56]. On the other hand, using finer bits for small values under log-quantization increases the communication complexity as compared to the uniform quantization [57]. This implies that log-quantization, in general, adds more complexity to the communication network as compared to uniform quantization. In terms of network structure in the presence of quantization, although exponential graphs are structured and can facilitate efficient quantized communication, their optimality gap might be considerably affected by quantization inaccuracy. In other words, the efficacy of the network structure might be dominated by the effect of quantization. These comments are investigated by the numerical experiments in Section IV.
@@ III-A Algorithm
This section presents our main algorithm for learning over networks. The methodology is based on two WB matrices $A_{\gamma}=[a^{\gamma}_{ij}]\in\mathbb{R}^{n\times n}$ and $B_{\gamma}=[b^{\gamma}_{ij}]\in\mathbb{R}^{n\times n}$ that capture the weights for data-sharing over the network, one for consensus on the states and the other for gradient-tracking dynamics. The index $\gamma:k\mapsto\Gamma$ denotes the associated network topology that may change over time, whereas $\Gamma$ denotes the set of all possible topologies of the data-sharing network. The graph is all-time connected, i.e., all the network topologies in $\Gamma$ are connected. The two matrices $A_{\gamma}$ and $B_{\gamma}$ can be designed independently, matrix $A_{\gamma}$ to characterize the weights on the local state value $\mathbf{x}_{i}$ and matrix $B_{\gamma}$ to characterize the weights on the auxiliary variable $\mathbf{y}_{i}$ for gradient tracking. These weight matrices are WB over $\mathcal{G}$ as discussed before, where the entries satisfy the following:
@@
The solution by (19)-(20) is composed of two dynamics: one to derive the agents to reach agreement on the optimization variable $\mathbf{x}$ and one to track the consensus on the local gradients. For the latter, we introduce an auxiliary variable $\mathbf{y}$ that tracks the average of gradients in a consensus-based setup and this variable derive the first dynamics toward the global minimum, while both dynamics address log-quantized communication. The log-scale quantization keeps the algorithm efficient in terms of communication with better performance as more resources are allocated near the origin. In contrast to most existing methods with weight-stochastic matrices (introduced in Section I), the proposed dynamics work under WB design and under switching network topologies. These make it easier to handle link-failure as discussed in Fig. 2. Moreover, the link (or channel) nonlinearity $h_{l}(\cdot)$ is not considered in the existing literature, which allows to address log-scale data quantization by setting $h_{l}(z)=q_{l}(z)$ given in Eq. (4). The discrete-time version of our methodology is summarized in Algorithm 1.
@@ III-B Proof of Convergence
This section proves the convergence of the distributed dynamics (19)-(20) using eigen-spectrum perturbation-based analysis. We also state some relevant lemmas. The proofs in this section hold for any sign-preserving, odd, sector-bound nonlinear mapping $h_{l}(\cdot)$, which also includes log-scale quantization $h_{l}(z)=q_{l}(z)$ defined by (4). In the rest of the paper, for notation simplicity, we drop the dependence $(t,\gamma)$ unless where it is needed.
@@
[59] Given the balanced digraph $\mathcal{G}$ its Laplacian matrix $\overline{A}$ (or $\overline{B}$) has all its eigenvalue in the left-half-plane except one isolated zero eigenvalue with associated non-negative left eigenvector $\mathbf{u}_{1}^{\top}$ satisfying $\sum_{i=1}^{n}u_{1,i}>0$.
@@
where $H:=\mbox{diag}[\nabla^{2}f_{i}(\mathbf{x})]$ and $M(\alpha)=M_{0}+\alpha M_{1}$ with
@@
For the rest of the proof, we study the eigenspectrum perturbation of $M^{0}$ and then extend the results to $M^{0}_{q}$ using Eq. (44). From Lemma 2, both Laplacian matrices  $\overline{A}$ and $\overline{B}$ have one isolated zero eigenvalue and the rest in the left-half-plane for all cases of the switching networks. Therefore, $m$ sets of (possibly changing) eigenvalues of $M_{0}$, associated with dimensions $j=\{1,\ldots,m\}$ are in the form,
@@
Next, we consider $\alpha M_{1}$ as a perturbation to matrix $M_{0}$. This is done using Lemma 1. In particular, we check how the perturbation term $\alpha M_{1}$ affects the two zero eigenvalues $\lambda_{1,j}$ and $\lambda_{2,j}$ of $M_{0}$. Denote the perturbed zero eigenvalues by $\lambda_{1,j}(\alpha)$ and $\lambda_{2,j}(\alpha)$ associated to $M$. For all WB and connected switching network topologies, the right and left unit eigenvectors of $\lambda_{1,j}$$\lambda_{2,j}$ follow from Lemma 2 and [59] as [^5] The normalizing factors of the unit vectors might be ignored as in the followings we only care about the sign of the terms in $U^{\top}M_{1}V$, not the exact values. ,
@@
where $u_{2,i}$ is the $i$th element of $\mathbf{u}_{2}$. Then, using Lemma 1, the perturbations  $\partial_{\alpha}d\lambda_{1,j}|_{\alpha=0}$ and $\partial_{\alpha}d\lambda_{2,j}|_{\alpha=0}$ are defined based on the eigenvalues of the matrix in (54). Since this matrix is triangular and from Eq. (55), $\partial_{\alpha}\lambda_{1,j}|_{\alpha=0}=0$ and $\partial_{\alpha}\lambda_{2,j}|_{\alpha=0}<0$. Therefore, for all sets of switching network Laplacians, the perturbation analysis imply that $\alpha M_{1}$ pushes $m$ zero eigenvalues $\lambda_{2,j}(\alpha)$ of $M$ toward the left-half-plane and the other $m$ zero eigenvalues $\lambda_{1,j}(\alpha)$ remain at zero. This implies that sufficiently small $\alpha$ gives the eigen-spectrum of $M$ as
@@
First, we determine the admissible bound on $\alpha$ for proof of convergence. Recall from [61, Appendix] that one can relate the spectrum of $M(\alpha)$ in (31) to $\alpha$. For ease of notation, from this point onward we derive the bound for $m=1$ (but it holds for any $m>1$). Performing row/column permutations in [61, Eq. (18)] $\sigma(M)$ can be determined from the following,
@@
Similar to the proof of Theorem 1, for stability we need to find the admissible range of $\alpha$ values for which the eigenvalues $\lambda$ remain in the left-half-plane, except one zero eigenvalue. The analysis here is based on the fact that the eigenvalues are continuous functions of the matrix elements [62]. It is clear that $\alpha=0$ satisfies Eq. (57) and gives $\mbox{det}((\overline{W}-\lambda I_{n})(\overline{A}-\lambda I_{n}))=0$. This leads to the eigen-spectrum following as $\sigma(M)=\sigma(\overline{A})\cup\sigma(\overline{W})$ with two zero eigenvalues for all switching network topologies. We need to find the other root $\overline{\alpha}>0$, which gives the admissible range as $0<\alpha<\overline{\alpha}$ for the stability of $M(\alpha)$. This follows the continuity of $\sigma(M)$ as a function of $\alpha$ [62]. With some abuse of notation, for any $\operatorname{Re}\{\lambda\}<0$ one can reformulate (57) as
@@
This implies that $\dot{\mathcal{V}}={\delta}^{\top}\dot{\delta}=\delta^{\top}M_{q}{\delta}$ and from [59, Sections VIII-IX], we have